The sum of two numbers is 22. Three times the smaller number minus two times the larger number is negitive 29. What is the smaller number?
Found 2 solutions by mananth, unlockmath:
Answer by mananth(16946)

let the numbers be x & y
x+y=22
3x-2y=-29
3(x+y)-(3x-2y)= 3*22-(-29)
3x+3y-3x+2y=66+29
5y=95
y= 19
x= 3
Ananth

Answer by unlockmath(1688)

Hello,
Let's have x & y represent the numbers. We can set up 2 equations as:
x+y=22
3x-2y=-29 Let's use substitution by changing the first equation to:
x=22-y
And plug this into the second.
3(22-y)-2y=-29 Rewritten as:
66-3y-2y=-29 Combine like terms and subtract 66 from both sides to get:
-5y=-95 Divide both sides by -5 to get:
y=19
Solve x by plugging this into either equation.
x+19=22
x=3
Check it out to see if these work.
RJ
www.math-unlock.com